# NaturalReasoning: Reasoning in the Wild with 2.8M Challenging Questions

### v2

@@ 5.1 Results
The scaling trends plotted by averaging performance on the three benchmarks are shown in Figure 6, and Table 2 provides a detailed breakdown of model performance across different dataset sizes and benchmarks.
@@
As shown in Figure 6, models trained on NaturalReasoning require fewer training examples to achieve superior performance. With just 1.5 million training examples, the model trained on NaturalReasoning already outperforms Llama3.1-8B-Instruct, which was extensively tuned for instruction-following with more data (Grattafiori et al., 2024). In contrast, other datasets, including OpenMathInstruct-2 and WebInstruct, fail to surpass Llama3.1-8B-Instruct even when trained on 2.8 million data points. This highlights that NaturalReasoning provides more useful reasoning supervision per training sample, making it a better choice for improving model reasoning abilities efficiently.
@@
A closer look at Table 2 reveals that OpenMathInstruct-2 consistently achieves the highest scores on the MATH benchmark, with performance increasing from 50.83 (500K) to 59.25 (2.8M). This confirms that OpenMathInstruct-2 is well-optimized for pure math reasoning. However, its performance on GPQA and MMLU-Pro is significantly weaker, where GPQA accuracy plateaus around 27–26 as dataset size increases, and MMLU-Pro accuracy fluctuates without significant improvement. This suggests that while OpenMathInstruct-2 provides strong supervision in math reasoning, it lacks the diversity required to generalize to broader scientific and multidisciplinary reasoning tasks.
@@
6 Reference Answer Usefulness
We demonstrate the potential usefulness of reference answers using questions from science-reasoning-15k. We remove the questions that we are not able to extract a reference answer for and conduct a comparison to understand the utility of reference answers. We fine-tune the Llama3.1-8B-Instruct model using data filtered by final answer verification and compare to a model trained on the unfiltered data.
| Training size | GPQA | MMLU | |
| Llama3.1-8B-Inst | – | 31.82 | 49.79 |
| Unfiltered SFT | 12349 | 31.82 | 49.92 |
| Filtered SFT | 7646 | 32.15 | 50.06 |
@@ 7 Unsupervised Self-Training
Since open-ended reasoning questions are difficult to evaluate using exact match with reference answers, we explore whether our dataset can facilitate self-training through either strong external reward models or self-reward mechanisms (Yuan et al., 2024a).
We use the science-reasoning-15k questions and sample 32 responses per question using the Llama3.1-8B-Instruct model. We evaluate different unsupervised self-training methods based on Rejection-based sampling Fine-Tuning (RFT) and Direct Preference Optimization (DPO) (Rafailov et al., 2023) across GPQA-Diamond and MMLU-Pro, focusing on the effectiveness of different reward scoring strategies. Each approach relies on sampling 32 candidate responses per question, followed by selecting responses based on reward scores. RFT employs rejection sampling, selecting the highest-scoring response for SFT training, while DPO constructs training pairs using both the highest and lowest-scoring responses. For external reward models, we consider Qwen2.5-Math-RM-72B (Yang et al., 2024) and INF-ORM-Llama3.1-70B222https://huggingface.co/infly/INF-ORM-Llama3.1-70B.
In addition to utilizing external reward models, we explore a self-rewarding framework where the model evaluates and assigns rewards to its own generated responses. Specifically, we consider the following self-rewarding strategies:
@@
Self-scoring: The model is prompted with both the question and a response and asked to assess whether the response is valid. We compute a reward score as the difference between the log probability of a “yes” judgement and a “no” judgement. The detailed prompt used for this evaluation is provided in Figure 11.
Self-scoring with filtering: on top of self-scoring, when applying RFT or DPO, we introduce an additional filtering mechanism to improve training quality. Specifically, for RFT, if the highest-ranked response has a self-score below zero, it is discarded. For DPO, if the preferred response in a pair has a self-score below zero, the pair is removed from training.
We train Llama3.1-8B-Instruct using RFT data and DPO data constructed through these four methods. The training setup includes a learning rate of , a batch size of 64, and training for three epochs, with checkpoints saved every 50 steps.
@@ 7.1 Results
| Model | GPQA | MMLU | AVG |
@@
| RFT training using external Reward model | |||
@@
| DPO training using external Reward model | |||
@@
While using external reward models, such as INF-ORM-Llama3.1-70B and Qwen2.5-Math-RM-72B, could outperform the baseline, self-reward methods achieve comparable or even superior results. Notably, self-score-filtered SFT and self-score-filtered DPO deliver the best performance on GPQA-Diamond (35.02), with self-score-filtered DPO achieving the highest overall score (43.67 AVG). These results highlight that self-reward mechanisms can effectively guide self-training without relying on external reward models.
@@ 8 Related Work
Synthetic data has emerged as a promising solution for improving performance and enabling new skills. Some approaches bootstrap new data from existing annotated data (e.g., STaR (Zelikman et al., 2022) augments with new CoT rationales and MetaMath (Yu et al., 2024a) rewrites the questions in MATH and GSM8K in several ways), but these techniques rely on the existence of a high-quality dataset. Other techniques such as that of OpenMathInstruct-2 (Toshniwal et al., 2024b), NuminaMath (Li et al., 2024b), Xwin-Math (Li et al., 2024a), and Self-Instruct (Wang et al., 2023) generate new data from only a few seed examples using an LLM but scaling to new domains or new content not previously seen in training remains a significant challenge. MMIQC (Liu et al., 2024a) parses QA pairs from Mathematics Stack Exchange, using the highest-ranked answer, but few measures are taken to curate for quality and the resulting dataset is also specific to the math domain. Similar to our work, WebInstruct (Yue et al., 2024) harvests question-answer pairs from pre-training corpora and spans multiple domains, but is dependent on carefully crafted rule-based filters.
@@
Most prior works typically depend on human-annotated (gold) final answers (Zelikman et al., 2022; Chen et al., 2024; Pang et al., 2024) or the use of an external reward model (Singh et al., 2024; Dong et al., 2023). However, manually annotating or verifying final answers is particularly resource-intensive for complex, multi-step problems and training effective reward models for reasoning often requires human evaluation of LLM outputs (Cobbe et al., 2021; Uesato et al., 2022; Lightman et al., 2023), making it similarly costly. Like works such as She et al. (2024); Yuan et al. (2024b); Rosset et al. (2024); Tran et al. (2023), our work explores self-training even in the absence of gold labels and does not limit itself to questions with short, easily verifiable targets.
@@ 9 Conclusion
Impact Statement
This paper seeks to improve reasoning capabilities of large language models through leveraging pretraining corpora. While our efforts are focused on curating high-quality, diverse data, models trained using this data may exhibit undesirable behavior not examined in our work. Therefore, comprehensive evaluation would be needed to evaluate and address any potential pre-existing or existing biases in LLMs which leverage this data.
@@ Appendix A Clustering Results
We present the results of our clustering in Figure LABEL:fig:topic_clusters. The procedure for producing this clustering is described in Section 3.3.
@@ C.1 Generation
Appendix D Prompts
The prompt we used for annotating reasoning from the document is shown in Figure 8. We additionally provide the prompt for annotating question validity and difficulty (Figure 9), the prompt used to check if a generated response matches the reference (Figure 10), and the prompt for self scoring (Figure 11).

### v3

@@ 5.1 Results
The scaling trends for Llama3.1-8B-Base model plotted by averaging performance on the three benchmarks are shown in 4(a), and 2(a) provides a detailed breakdown of model performance across different dataset sizes and benchmarks. The scaling trends for Qwen2.5-7B model also show the superiority of NaturalReasoning and we put the results in Appendix E.
@@
As shown in 4(a), models trained on NaturalReasoning require fewer training examples to achieve superior performance. With just 1.5 million training examples, the model trained on NaturalReasoning already outperforms Llama3.1-8B-Instruct, which was extensively tuned for instruction-following with more data (Grattafiori et al., 2024). In contrast, other datasets, including OpenMathInstruct-2 and WebInstruct, fail to surpass Llama3.1-8B-Instruct even when trained on 2.8 million data. Each NaturalReasoning sample therefore provides denser, more effective reasoning supervision, making it the most data-efficient choice for boosting model reasoning performance.
@@
A closer look at 2(a) reveals that OpenMathInstruct-2 consistently achieves the highest scores on the MATH benchmark, with performance increasing from 50.83 (500K) to 59.25 (2.8M). This confirms that OpenMathInstruct-2 is well-optimized for pure math reasoning. However, its performance on GPQA and MMLU-Pro is significantly weaker, where GPQA accuracy plateaus around 27–26 as dataset size increases, and MMLU-Pro accuracy fluctuates without significant improvement. This suggests that while OpenMathInstruct-2 provides strong supervision in math reasoning, it lacks the diversity required to generalize to broader scientific reasoning tasks.
@@
6 Eliciting Long Chain-of-Thought
In addition to the emergence of OpenAI-o1 and DeepSeek-R1, several studies suggest that simpler tasks require fewer steps while complex tasks benefit significantly from longer CoTs (Jin et al., 2024). When encouraging the model to think for longer, they are able to solve questions that they previously could not. Motivated by this, we investigate whether the questions in NaturalReasoning are complex enough to benefit from longer CoTs from a stronger reasoning model. We do so by distilling Deepseek-R1 responses to Llama3.3-70B-Instruct.
We randomly sampled 1K questions from NaturalReasoning and used SGLang (Zheng et al., 2023) to prompt DeepSeek-R1, generating one response per question. Resulting response lengths range from 745 to 14.6K tokens with an average length of 4430 tokens. We then supervised finetune Llama-3.3-70B-Instruct on this set and compare its performance against training on two strong, heavily curated datasets: s1K-1.1 (Muennighoff et al., 2025) and LIMO (Ye et al., 2025). Both datasets underwent multiple filtering stages to ensure that their questions are high-quality, diverse, and challenging; for consistency, all responses were generated with DeepSeek-R1. To keep the evaluation consistent with the setting used in Guo et al. (2025), we report pass@1 averaged across samples. Each sample is generated using temperature=, top_p=.
The results in Table 2 show that a randomly selected subset of 1k questions from NaturalReasoning matches—even slightly exceeds—the performance obtained on datasets that underwent several rounds of meticulous filtering and curation. This parity underscores that NaturalReasoning contains questions that are diverse, challenging, and of consistently high quality.
To examine the impact of scale, we expanded the random sample size from 1K to 10K and finally to 100K NaturalReasoning questions; the corresponding results are also presented in Table 2. Performance increases monotonically across every benchmark, providing further evidence that enlarging the slice of NaturalReasoning delivers substantial gains precisely because the added questions maintain the same high standard of quality. Moreover, fine-tuning Llama-3.3-70B-Instruct on just 100K randomly sampled questions from NaturalReasoning brings the model close to DeepSeek-R1-Distill-Llama-70B, which was trained on 800K examples. Despite using only one-eighth the data, our model outperforms DeepSeek-R1-Distill-Llama-70B on GPQA-Diamond and MMLU-Pro, and falls only slightly behind on MATH-500. This further attests to both the scalability and the intrinsic quality of NaturalReasoning.
| Training size | GPQA-Diamond | MMLU-Pro | MATH-500 | Average | |
| Llama3.3-70B-Instruct | 0 | 50.5 | 70.5 | 77.0 | 66.0 |
| LIMO | 817 | 56.5 | 76.8 | 86.6 | 73.3 |
| s1K-1.1 | 1,000 | 62.7 | 77.4 | 86.6 | 75.6 |
| NaturalReasoning | 1,000 | 63.5 | 78.0 | 86.2 | 75.9 |
| NaturalReasoning | 10,000 | 65.6 | 78.4 | 87.4 | 77.1 |
| NaturalReasoning | 100,000 | 67.3 | 79.5 | 89.8 | 78.9 |
| DeepSeek-R1-Distill-Llama-70B | 800,000 | 65.2 | 78.5 | 94.5 | 79.4 |
@@ 7 Unsupervised Self-Training
Since open-ended reasoning questions are difficult to evaluate using exact match with reference answers, we explore whether our dataset can facilitate self-training through either strong external reward models or self-reward mechanisms (Yuan et al., )
To test the effectiveness of self-training without confounding factors from distribution shift, we evaluate on GPQA-Diamond as test set, and use the remaining questions from GPQA as seeds to retrieve similar questions from NaturalReasoning. We curated 15,000 questions in total, which we refer to as SelfTrain-15k. More details are in Section C.2. Models trained on this subset can still be evaluated on the GPQA-Diamond test set as those questions are not used for data selection.
We verify unsupervised self-training under two different training method: Rejection-based sampling Fine-Tuning (RFT) and Direct Preference Optimization (DPO) (Rafailov et al., 2023), focusing on the effectiveness of different reward scoring strategies. Each approach relies on sampling 32 candidate responses per question, followed by selecting responses based on reward scores. RFT employs rejection sampling, selecting the highest-scoring response for SFT training, while DPO constructs training pairs using both the highest and lowest-scoring responses. For external reward models, we consider Qwen2.5-Math-RM-72B (Yang et al., 2024) and INF-ORM-Llama3.1-70B333https://huggingface.co/infly/INF-ORM-Llama3.1-70B. In addition, we explore a self-rewarding framework where the model evaluates and assigns rewards to its own generated responses. Specifically, we consider the following self-rewarding strategies:
@@
Self-scoring: The model receives the question and candidate response in a single prompt and is asked to assess whether the response is valid. We define the reward as the log-probability difference between the judgements “yes” and “no”. The full prompt is in Figure 11.
Self-scoring with filtering: on top of self-scoring, when applying RFT or DPO, we introduce an additional filtering mechanism. Specifically, for RFT, if the highest-ranked response has a self-score below zero, it is discarded. For DPO, if the preferred response in a pair has a self-score below zero, the pair is removed from training.
We train Llama3.1-8B-Instruct using RFT data and DPO data constructed through these methods. We use learning rate of , batch size of 64, and train for three epochs, with checkpoints saved every 50 steps. We report test performance on GPQA-Diamond and MMLU-Pro in Table 3.
@@ 7.1 Results
| Model | GPQA-Diamond | MMLU-Pro | Average |
@@
| RFT training using external reward model & self-reward | |||
@@
| DPO training using external reward model & self-reward | |||
@@
While using external reward models, such as INF-ORM-Llama3.1-70B, could outperform the baseline, self-reward methods achieve comparable or even superior results. Notably, self-score-filtered SFT and self-score-filtered DPO deliver the best performance on GPQA-Diamond (35.02), with self-score-filtered DPO achieving the highest overall score (43.67). These results highlight that self-reward mechanisms can effectively guide self-training without relying on external reward models.
@@ 8 Related Work
Synthetic data has emerged as a promising solution for improving performance. Some approaches bootstrap new data from existing data (e.g., STaR (Zelikman et al., 2022) augments with new CoT rationales and MetaMath (Yu et al., ) rewrites the questions in MATH and GSM8K in several ways), but these techniques rely on the existence of a high-quality dataset. Other techniques such as that of OpenMathInstruct-2 (Toshniwal et al., ), Xwin-Math (Li et al., 2024a), and Self-Instruct (Wang et al., 2023) generate new data from only a few seed examples using an LLM but scaling to new domains remains a significant challenge. MMIQC (Liu et al., ) parses QA pairs from Mathematics Stack Exchange, using the highest-ranked answer, but few measures are taken to curate for quality and the resulting dataset is also specific to the math domain. Similar to our work, WebInstruct (Yue et al., 2024) harvests question-answer pairs from pre-training corpora and spans multiple domains, but is dependent on carefully crafted rule-based filters.
@@
Most prior works typically depend on human-annotated (gold) final answers (Zelikman et al., 2022; Chen et al., 2024; Pang et al., 2024) or the use of an external reward model (Singh et al., 2024; Dong et al., 2023). However, manually annotating or verifying final answers is particularly resource-intensive for complex, multi-step problems and training effective reward models for reasoning often requires human evaluation of LLM outputs (Cobbe et al., 2021; Uesato et al., 2022; Lightman et al., 2023), making it similarly costly. Like works such as She et al. (2024); Yuan et al. (2024); Rosset et al. (2024); Tran et al. (2023), our work explores self-training in the absence of gold labels and does not limit itself to questions with short, easily verifiable targets.
@@ 9 Conclusion
Limitation & Impact Statement
Although our study already validates NaturalReasoning ’s value for large-scale offline training—covering supervised distillation and preference-based self-training (RFT, DPO)—we did not pursue complementary directions such as online RL with reference-answer rewards or an exhaustive survey of reward models; these remain natural extensions for future work. This paper seeks to improve reasoning capabilities of large language models through leveraging pretraining corpora. While our efforts are focused on curating high-quality, diverse data, models trained using this data may exhibit undesirable behavior not examined in our work. Therefore, comprehensive evaluation would be needed to evaluate and address any potential pre-existing or existing biases in LLMs which leverage this data.
@@ Appendix A Clustering Results
We present the results of our clustering in Figure 5. The procedure for producing this clustering is described in Section 3.3.
@@ C.1 Generation
C.2 SelfTrain-15k Curation
Specifically, for each question in the non-Diamond subset of GPQA, we retrieve the 1,000 most similar questions from NaturalReasoning, which were already decontaminated against the entire GPQA datset. Similarity is computed using cosine similarity between two question embeddings. After obtaining this candidate set, we apply deduplication and perform clustering, grouping the questions into 15,000 clusters. From each cluster, we select the questions closest to the cluster center, ensuring a diverse and representative dataset for downstream science reasoning tasks. This process resulted in a pool of 15,000 questions, which we refer to as SelfTrain-15k.
Appendix D Human Evaluation of Question Quality
To assess question quality reliably, we also conducted a human evaluation by randomly sampling 100 questions from each dataset. Two expert annotators independently rated each question, and we used the average of their scores as the final quality measure. The results, shown in Table 5, confirm that the NaturalReasoning dataset consistently produces higher-quality questions.
| Dataset | MetaMathQA | OpenMathInstruct-2 | NuminaMath | WebInstruct | NaturalReasoning |
| Quality score | 5.09 | 5.37 | 5.61 | 5.92 | 6.46 |
Appendix E Qwen2.5-7B Scaling Results
The scaling trends for the Qwen2.5-7B model plotted by averaging performance on the three benchmarks are shown in 6(a). 6(a) provides a detailed breakdown of model performance across different dataset sizes and benchmarks. It is clear that NaturalReasoning shows superior scaling trends than other datasets.
| SFT Dataset | 500K | 1M | 1.5M |
| MATH | |||
| WebInstruct | 38.82 | 40.12 | 38.49 |
| NuminaMath | 62.68 | – | – |
| OpenMathInstruct-2 | 64.25 | 65.51 | 65.55 |
| NaturalReasoning | 65.85 | 66.26 | 66.67 |
| GPQA | |||
| WebInstruct | 31.85 | 30.28 | 29.76 |
| NuminaMath | 1.71 | – | – |
| OpenMathInstruct-2 | 18.23 | 14.21 | 9.08 |
| NaturalReasoning | 16.44 | 20.76 | 21.35 |
| MMLU-Pro | |||
| WebInstruct | 43.65 | 42.77 | 43.18 |
| NuminaMath | 51.18 | – | – |
| OpenMathInstruct-2 | 48.96 | 48.88 | 49.36 |
| NaturalReasoning | 51.37 | 52.96 | 53.10 |
Appendix F Reference Answer Usefulness
We demonstrate the potential usefulness of reference answers using questions from SelfTrain-15k. We remove the questions that we are not able to extract a reference answer for and conduct a comparison to understand the utility of reference answers. We fine-tune the Llama3.1-8B-Instruct model using data filtered by final answer verification against a model trained on the unfiltered data.
| Training size | GPQA-Diamond | MMLU-Pro | Average | |
| Llama3.1-8B-Instuct | – | 31.82 | 49.79 | 40.81 |
| Unfiltered SFT | 12349 | 31.82 | 49.92 | 40.87 |
| Filtered SFT | 7646 | 32.15 | 50.06 | 41.11 |
Appendix G Prompts
The prompt we used for annotating reasoning from the document is shown in Figure 7, Figure 8. We additionally provide the prompt for annotating question validity and difficulty (Figure 9), the prompt used to check if a generated response matches the reference (Figure 10), and the prompt for self scoring (Figure 11).